Help me solve g(x)=absolute value of x -3
Answer by venugopalramana(3286)

IN ABSOLUTE VALUE PROBLEMS YOU HAVE TO CONSIDER 2 CASES
CASE 1..THE GIVEN EXPRESSION IS POSITIVE...IN SUCH A CASE WE GET HERE
X-3 IS POSITIVE OR X>3....THEN G(X)=+(X-3)=X-3........I
CASE 2..THE GIVEN EXPRESSION IS NEGATIVE...IN SUCH A CASE WE GET HERE
X-3 IS NEGATIVE OR X<3....THEN G(X)=-(X-3)= -X+3.........II
THESE 2 EQNS I AND II ARE CALLED THE SOLUTION SET WITH THE ASSUMED PARAMETERS ON X BEING MENTIONED AS CONDITIONS...THAT IS...
IF X>3...THEN...G(X)=X-3...................I
IF X<3...THEN....G(X)= -X+3